# Artificial Intelligence in Creative Industries: Advances Prior to 2025

### v2

@@ 3.6 Data Compression
Data compression plays an important role in various applications of creative technologies, which effectively reduces memory space and bandwidth requirements during signal storage and transmission [287]. Although standard coding methods based on conventional signal processing theories are still widely employed in most application scenarios, learning-based solutions have emerged in research, showing great potential to achieve competitive performance in recent years. This subsection provides a brief overview on the recent advances in image, video and audio compression, in particular focusing on the approaches proposed after 2021.
@@ 3.6.1 Image Compression
Since the first neural image codec [288] was proposed in 2016, numerous learning-based image compression methods have been developed, with significant performance improvements reported [289, 290]. In recent years, inspired by the latest advances in neural network architectures, many neural image codecs have been proposed, which outperform the latest standard image codecs. Instead of using CNNs as the basic network structures, transformer-based architectures have become popular, with potential to achieve better compression efficiency. Notable examples include SwinT-ChARM [201], STF [202] and LIC-TCM [203]. SwinT-ChARM [201] employs Swin transformers for non-linear transforms and outperforms the latest standard image codec, the Versatile Video Coding (VVC) Test Model (VTM, All Intra). STF [202] is based on a symmetrical transformer framework containing absolute transformer blocks in both the down-sampling encoder and the up-sampling decoder, which also shows improved rate quality performance over VTM. LIC-TCM [203] exploits the local modeling ability of CNN and the non-local modeling performance of transformers, and proposes a parallel transformer-CNN mixture block. This new network structure, together with a channel-wise entropy model based on attention modules using Swin transformers, contributes to the superior performance of STF, with a more than 10% bitrate saving over VTM.
@@
There is another group of learned image coding methods, which are based on advanced generative models. Early works [291, 292] employ GANs to generate more photo realistic results with improved visual quality. Although these models fail to outperform conventional, CNN-based or transformer-based approaches, when distortion-based quality metrics, e.g., PSNR, are used for performance evaluation, they have been reported to perform well when perceptual quality models, such as MS-SSIM [293] and VMAF [294], or subjective tests are employed to measure perceived video quality. More recently, diffusion models have been applied in image compression to allow realistic reconstruction at ultra-low bitrates [204] and achieve more competitive performance over GAN-based models [205]. However, it should be noted that some of these generative models aim to generate (or synthesize) images with “perfect realism” rather than reconstruct results which are similar to the original content. Notable works in this category also include the image codec using score-based generative models [206] and the diffusion-based residual augmentation codec (DIRAC) [207]. Moreover, as another type of generative models, INR has been employed for image compression, which learns a mapping between the spatial coordinates and the respective pixel values for the input image. The learned INR model is then compressed through parameter quantization and model compression to minimize the required bitrate. Notable INR-based image codecs including COIN/COIN++ [209, 210] and [211] that combines SIREN networks [208] with positional encoding,
@@
In order to evaluate and compare neural image codecs under fair test conditions, public grand challenges have been held by different organizers, which are typically associated with international conferences. One of the most well-known ones is the Challenge on Learned Image Compression (CLIC) [295], and in its latest competition, the best performing learned image codec offers up to 0.6dB PSNR gain over VTM (version 22.2, All Intra) at similar bitrates, which is based on an autoencoder architecture with latent refinement and perceptual losses.
@@
To support the deployment of neural image codecs, the International Organization for Standardization (ISO)/International Electrotechnical Commission(IEC) have jointly started to develop a royalty-free learned image coding standard, denoted as JPEG AI [296], which aims to offer significant performance improvement over existing standards for both human and machine vision tasks. The Call for Proposals of JPEG AI was published in 2022, while the Working Draft and the Committee Draft outlining its core coding system were released in 2023 [297], with its first version published in October 2024 [297]. JPEG AI follows the same framework (the auto-encoder structure) as most existing neural image codecs, and its test model JPEG AI VM (version 4.3) has been reported to achieve up to 28.5% coding gains over VVC VTM (All Intra mode) [298].
@@ 3.6.2 Video Compression
Compared to images, the compression of video content is a much more challenging task, in particular with various immersive video formats and diverse content. Although video coding standards including H.264/AVC (Advanced Video Coding), H.265/HEVC (High Efficiency Video Coding) and H.266/VVC (Versatile Video Coding) are still predominant in real-world applications, learning-based video coding has been extensively investigated in the past five years, with deep learning enhanced conventional coding tools and end-to-end optimized neural video coding frameworks proposed.
@@
The enhancement of conventional coding tools focuses on employing deep learning techniques to improve the performance of a (or multiple) coding modules in a standard-applicant codec, including intra prediction [299], inter prediction [300], in-loop filtering [301], post filtering [302] and resolution re-sampling [303]. To facilitate efficient integration, the MPEG Joint Video Experts Team (JVET) has built a test model in 2022 based on VTM 11, named Neural Network-based Video Coding (NNVC) [304], with its latest version NNVC-7.1 containing two major learning-based coding tools, neural-network based intra prediction and in-loop filtering, which has achieved an up to 13% coding gain over VTM 11 (Random Access mode) [305]. However, this learning-based codec requires much higher computational complexity (up to 477 kMACs/pixel) and high-spec GPU support compared to conventional codecs. Meanwhile, members of the Alliance of Open Media (AOM) have also developed multiple CNN-based coding tools for the next generation of video coding standard beyond AV1. The latest proposals focus on the trade-off between performance and complexity, with one of them based on inloop filtering and super-resolution, which achieves an average BD-rate saving of 3.9% (in PSNR) over AVM, the test model of AV2, but only requires a much lower computational complexity (below 1.5kMACs/pixel) [306]. More recently, research has been conducted to further improve the performance of these learning-based coding tools utilizing more advanced network architectures, including ViTs [307], and diffusion models [214]. There are also investigations on applying preprocessing before compression [308, 309], where the training of the deep preprocessors is based on proxy video codecs and/or rate-distortion loss functions to simulate the behavior of conventional video coding algorithms.
@@
Alongside enhancing coding tools in conventional video codecs, more research activities investigate using neural networks to implement the whole coding workflow, which enables data-driven end-to-end optimization. The performance of these neural video codecs has improved significantly in the last five years, since the first attempt, DVC [310], was published, which only matches the performance of a fast implementation of H.264 (x264). When this paper was written, some latest well-performing learned video coding algorithms (e.g., DCVC-FM [311] and DCVC-LCG [312]) have already been able to compete or even outperform the state-of-the-art standard codecs, such as VVC VTM under certain coding configurations. These learning-based methods often focus on enhancement from different perspectives, including feature space conditional coding (e.g., FVC [313] and DCVC [314]), instance adaptation  [315, 316], and motion estimation (e.g., DCVC-DC [317]). New architectures have also been proposed such as CANF-VC [318] based on a video generative model, MTMT [212] using a masked image modeling transformer based entropy model and VCT [213] based on a video compression transformer. It is noted that although promising coding performance has been achieved in the aforementioned works, these neural video codecs (in particular those based on autoencoder backbones) are typically associated with high computational complexity (especially in the decoder), which constrains their deployment for practical applications. To address this issue, researchers attempted to achieve complexity reduction while maintaining the coding performance through model pruning and knowledge distillation [319, 320].
@@
It should be noted that the neural codecs mentioned above are typically trained offline with diverse video content [321], and deployed online for inference. In this case, model generalization becomes important, and this is why these codecs often have a large model capacity, resulting in large model sizes and slow inference runtime. Inspired by recent advances in implicit neural representations (INR), a new type of video codec has emerged that employs INR models to “represent” the video to be encoded by learning a coordinate-based mapping and compressing the network parameters for transmission. This approach converts a video coding problem into a model compression task, which allows the use of a much smaller network to “overfit” the input video, with the real potential for fast decoding. Existing implicit neural video representation (NeRV) models can be classified into index-based and content-based methods. The former takes frame [215], patch [216] or disentangled spatial/grid coordinates [322] as model input, while content-based approaches [42, 217, 218] have content-specific embedding as inputs. Currently, one of the best INR-based video codecs [219] has already achieved a performance similar to that of VVC VTM (RA), but with a much lower decoding complexity compared to autoencoder-based neural codecs. Some of these models have also been applied to volumetric video content [221, 222], demonstrating their great potential to compete with standard and other learning-based methods. However, it should be noted that the training of most NeRV models is based on entire video sequences or even datasets, which results in a high system delay and does not meet the requirement of many video streaming applications. To address this limitation, efforts have been made [220] towards more practical INR-based video compression (such as the Low Delay and Random Access modes in VVC VTM [323]) by combining pre-training and online model overfitting.
@@
Similarly to image compression, there have been grand challenges organized in recent years for neural video compression, with notable venues including the NN-based Video Coding Grand Challenge associated with The IEEE International Symposium on Circuits and Systems (ISCAS) and the Challenge on Learned Image Compression (CLIC, video coding track) with IEEE/CVF CVPR and Data Compression Conference (in 2024). The best performer in ISCAS 2024 NN-based Video Coding Grand Challenge Challenge offers an overall 55% BD-rate saving over HEVC Test Model HM, while the winner of the CLIC (video coding track) in 2024, a neural-network enhanced ECM codec, shows a more than 2dB (in PSNR) gain compared to VTM (RA) at the same bitrates.
@@ 3.6.3 Audio Compression
Similarly to images and videos, learning-based solutions have also been researched to compress audio signals, and most neural audio codecs are based on VQ-VAE [240]. SoundStream [324] is one of such models, which can encode audio content at various bitrate. It is based on a residual vector quantizer (RVQ) which trades off between rate, distortion, and complexity. This work has been further enhanced with a multi-scale spectrogram adversary and a loss balancer mechanism, resulting in improved rate-distortion performance. A more advanced universal model has been further developed [325] based on improved adversarial and reconstruction losses, which can compress different types of audio. RVQ has also been extended from a single scale to multiple scales [326], which performs hierarchical quantization at variable frame rates.
More recently, researchers have started to exploit the use of LLMs for audio compression, leveraging the audio generation/synthesis abilities of generative models. UniAudio 1.5 [327] is one of such attempts, which converts an audio into the textural space, which can be represented by a pre-trained LLM that shares a similar backbone of UniAudio [328], an universal audio foundation model. LFSC is another neural audio codec based on LLMs, which achieved fast LLM training and inference through finite scalar quantization and adversarial training.
@@ 3.7 Visual Quality Assessment
Assessing the quality of visual signals remains an important and challenging task for many image and video processing applications. While subjective tests involving human participants offer a golden rule, objective quality models are designed and used in many cases because of their time and cost efficiency. These quality assessment methods are typically used to effectively evaluate the performance of different visual processing approaches, and they can also be converted to loss functions, which are employed for optimizing learning-based processing models.
In recent years, quality assessment methods have also been enhanced by deep learning techniques. The resulting learning-based quality models can quickly adapt to a specific type of content, leading to better performance compared to conventional, hand-crafted quality metrics. This section provides a brief summary of existing works in this research area, and highlights the main challenges which should be addressed in the near future. A more comprehensive overview of the image and video quality assessment literature can be found in [329, 330, 331].
@@ 3.7.1 Quality assessment models
Image and video quality assessment methods can be classified into two primary categories according to the availability of the corresponding reference content to the distorted test version: full-reference and no-reference models [^47] Reduced-reference quality metrics do existing in the literature, but the research in this field is less active in recent years.. Prior to the AI era, conventional visual quality methods often exploit different characteristics of the human vision system and capture relevant information related to structural similarity (such as in SSIM and its variants [332, 333, 334]), distortion [335, 336, 337], and artifacts [338, 339, 340]. In many cases, the extract features are further processed by models that simulate texture masking [341], contract sensitivity [342], and saliency [343]. These hand-crafted quality models have also been combined with features within a regression-based framework in order to achieve more accurate prediction performance - VMAF is one of such examples [294]. When neural networks are involved for feature extraction, they are trained to capture information which can directly contribute to quality prediction through an end-to-end optimization strategy. Initially, convolution network networks were typically used, with notable examples such as DeepQA [344], LPIPS [345] and CONTRIQUE [346] for image quality assessment, and TLVQA [347], C3DVQA [348] and DeepVQA [349] for video quality assessment. Recent works have been reported to achieve better performance when Vision Transformers (ViTs) (or similar variants) are employed due to the effectiveness of their self-attention mechanism. Important works in this class including IQT [223], TRes [224], SaTQA [225], FastVQA [226] and RankDVQA [227]. The former has been further extended as DOVER [228] and COVER [229] when aesthetic and/or semantic aspects in the content are taken into account.
@@
More recently, inspired by the success of large language models (LLMs) [4, 350] in other machine learning tasks, they have also been utilized in image and video quality assessment and demonstrate great potential to achieve better model generalization. Q-Bench [351] is one of the first attempts that employs multimodality large language models to predict the perceptual quality of images based on prompt-driven evaluation. It inquiries the LLMs to provide information related to the final quality rating of the input image and the quality description. This has been further extended for video quality assessment tasks, named Q-Align [352]. Other notable works of this kind include X-iqe [353] that performs the quality prompt in a multi-iteration manner focusing on both image fidelity and aesthetics. Prompt-based approaches have also been proposed for differentiating the quality difference between multiple images, such as 2AFC-LMMs [354] based on a two-alternative forced choice prompt and MAP (maximum a posteriori) estimation. Moreover, recent research works also focus on using pre-trained vision-language models, such as CLIP [238], which align better image and text modalities. Important examples in this class for image quality assessment including ZEN-IQA [355], QA-CLIP [356] and PromptIQA [357]. Similar works have also been proposed for video quality assessment, such as BVQI [358, 359] and COVER [229].
@@
To support the training and validation of learning-based quality assessment models, image or video databases containing ground-truth subjective quality scores are typically employed. Commonly used image quality databases include LIVE [360], CSIQ [336], TID2013 [361], PieAPP and PIPAL, while video quality databases such as LIVE-VQA [362], KoNViD-1K [363], YouTube UGC [364] and LIVE-VQC [365] are typically employed for benchmarking in the literature. There are also databases developed that investigate the impact of specific video formats and/or artifacts, such as LIVE-YT-HFR [366] focusing on frame rates, VSR-QAD [367] on spatial resolution (or super-resolution artifacts), BAND-2k [368] on banding artifacts and Maxwell [359]/BVI-Artifact [369] containing multiple artifacts commonly produced in video streaming. Based on these databases, many learning-based quality assessment models are trained to minimize the difference (L1 or L2 norm) between predicted quality indices and subjective scores. However, due to the limited number of ground-truth quality labels associated with these databases and the resourcing costing nature for collecting subjective data through human participants involved psychophysical experiments, this type of training methodology cannot offer satisfactory performance, in particular when the model capacity is large. Moreover, since the experimental settings and conditions used for quality labeling are different in these databases, intra-database cross-validation is always required due to the limited model generalization and potential overfitting problems.

### v3

@@ 3.6 Data Compression
Data compression plays an important role in the delivery of creative content to audiences, effectively reducing memory and bandwidth requirements during signal storage and transmission [328]. Although coding methods based on conventional signal processing theories are still widely employed in most standards and application scenarios, learning-based solutions have emerged in research, showing great potential to achieve competitive performance in recent years. This subsection provides a brief overview of the recent advances in image, video, and audio compression, in particular focusing on the approaches proposed after 2021.
@@ 3.6.1 Image Compression
Since the first neural image codec [329] was proposed in 2016, numerous learning-based image compression methods have been developed, with significant performance improvements reported [330, 331]. Driven by the latest advances in neural network architectures, neural image codecs now outperform standard image codecs. Instead of using CNNs as the basic network structure, transformer-based architectures have become popular, offering the potential for better compression efficiency. Notable examples include SwinT-ChARM [226], STF [227] and LIC-TCM [228]. SwinT-ChARM [226] employs Swin transformers for non-linear transforms and outperforms the latest standard image codec, the Versatile Video Coding (VVC) Test Model (VTM, All Intra). STF [227] is based on a symmetrical transformer framework containing absolute transformer blocks in both the down-sampling encoder and the up-sampling decoder, which also shows improved rate-quality performance over VTM. LIC-TCM [228] exploits the local modeling ability of CNN and the non-local modeling performance of transformers, and proposes a parallel transformer-CNN mixture block. This new network structure, together with a channel-wise entropy model based on attention modules using Swin transformers, contributes to the superior performance of STF, with a more than 10% bitrate saving over VTM.
@@
An alternative approach to learned image coding is based on advanced generative models. Early works [332, 333] employed GANs to generate more photo realistic results with improved visual quality. Although these models fail to outperform conventional, CNN-based or transformer-based approaches, when distortion-based quality metrics, e.g., PSNR, are used for performance evaluation, they have been reported to perform well when perceptual quality models, such as MS-SSIM [334] and VMAF [335], or subjective tests are employed to measure perceived video quality. More recently, diffusion models have been applied in image compression to allow realistic reconstruction at ultra-low bitrates [229] achieving competitive performance compared to GAN-based models [230]. However, it should be noted that some of these generative models aim to generate (or synthesize) images with “perfect realism” rather than reconstruct results which are most similar to the original content. Notable work in this category includes image codecs using score-based generative models [231] and the diffusion-based residual augmentation codec (DIRAC) [232]. Moreover, another type of generative model based on INR has been employed for image compression; this learns a mapping between the spatial coordinates and the respective pixel values for the input image. The learned INR model is then compressed through parameter quantization and model compression to minimize the required bitrate. Notable INR-based image codecs include COIN/COIN++ [234, 235] and [236] that combine SIREN networks [233] with positional encoding.
@@
In order to evaluate and compare neural image codecs under fair test conditions, public grand challenges have been increasingly run, typically associated with international conferences. One of the most well-known of these is the Challenge on Learned Image Compression (CLIC) [336]. In its latest competition, the best performing learned image codec [337], which is based on a GAN-enhanced Vector Quantized Variational AutoEncoder (VQ-VAE) framework, offered up to 0.6dB PSNR gain over VTM (version 22.2, All Intra) at similar bitrates; this codec is based on an autoencoder architecture with latent refinement and perceptual losses.
@@
To support the deployment of neural image codecs, the International Organization for Standardization (ISO)/International Electrotechnical Commission(IEC) has developed a royalty-free learned image coding standard, denoted as JPEG AI [338], which aims to offer significant performance improvement over existing standards for both human and machine vision tasks. The Call for Proposals of JPEG AI was published in 2022, while the Working Draft and the Committee Draft outlining its core coding system were released in 2023 [339], with its first version published in October 2024 [339]. JPEG AI follows the same framework (the auto-encoder structure) as most existing neural image codecs, and its test model JPEG AI VM (version 4.3) has been reported to achieve up to 28.5% coding gains over VVC VTM (All Intra mode) [340].
@@ 3.6.2 Video Compression
Compared to image coding, the compression of video content is a much more challenging task, particularly for immersive video formats and diverse content types. Although video coding standards including H.264/AVC (Advanced Video Coding), H.265/HEVC (High Efficiency Video Coding) and H.266/VVC (Versatile Video Coding) are still predominant in real-world applications, learning-based video coding has advanced dramatically in the past five years, with new deep learning enhanced conventional coding tools and end-to-end optimized neural video coding frameworks proposed.
@@
i) The enhancement of conventional coding tools focuses on employing deep learning techniques to improve the performance of one (or multiple) coding modules in a standard-applicant codec. These modules include intra prediction [341], inter prediction [342], in-loop filtering [343], post filtering [344] and resolution re-sampling [345]. To facilitate efficient integration, the MPEG Joint Video Experts Team (JVET) built a test model in 2022 based on VTM 11, named Neural Network-based Video Coding (NNVC) [346], with its latest version NNVC-7.1 containing two major learning-based coding tools, neural-network based intra prediction and in-loop filtering, which has achieved an up to 13% coding gain over VTM 11 (Random Access mode) [347]. However, this learning-based codec requires much higher computational complexity (up to 477 kMACs/pixel) and high-spec GPU support compared to conventional codecs. Meanwhile, members of the Alliance of Open Media (AOM) have also developed multiple CNN-based coding tools for the next generation of video coding standard beyond AV1. The latest proposals focus on the trade-off between performance and complexity, with one of them based on inloop filtering and super-resolution, which achieves an average BD-rate saving of 3.9% (in PSNR) over AVM, the test model of AV2, but only requires a much lower computational complexity (below 1.5kMACs/pixel) [348]. More recently, research has been conducted to further improve the performance of these learning-based coding tools utilizing more advanced network architectures, including ViTs [349], and diffusion models [239]. There are also investigations on applying preprocessing before compression [350, 351], where the training of the deep preprocessors is based on proxy video codecs and/or rate-distortion loss functions to simulate the behavior of conventional video coding algorithms.
@@
ii) End-to-end optimized neural video codecs. Alongside the enhancement of coding tools in conventional video codecs, more recent research activities have focused on using neural networks to implement the whole coding workflow, enabling data-driven end-to-end optimization. The performance of these neural video codecs has advanced significantly in the last five years, since the first attempt, DVC [352], was published. DVC matched the performance of a fast implementation of H.264 (x264). However currently, learned video coding algorithms (e.g., DCVC-FM [353] and DCVC-LCG [354]) are able to compete or even outperform the state-of-the-art standard codecs, such as VVC VTM under certain coding configurations. These learning-based methods often focus on enhancement from different perspectives, including feature space conditional coding (e.g., FVC [355] and DCVC [356]), instance adaptation  [357, 358], and motion estimation (e.g., DCVC-DC [359]). New architectures have also been proposed such as CANF-VC [360] based on a video generative model, MTMT [237] using a masked image modeling transformer-based entropy model and VCT [238] based on a video compression transformer. It is noted that although promising coding performance has been achieved in the aforementioned works, these neural video codecs (in particular those based on autoencoder backbones) are typically associated with high computational complexity (especially in the decoder), which constrains their deployment for practical applications. To address this issue, researchers attempted to achieve complexity reduction while maintaining the coding performance through model pruning and knowledge distillation [361, 362].
@@
It should be noted that the neural codecs mentioned above are typically trained offline with diverse video content [363], and deployed online for inference. In this case, model generalization becomes important, and this is why these codecs often have a large model capacity, resulting in large model sizes and slow inference runtime. Inspired by recent advances in implicit neural representations (INR), a new type of video codec has emerged that employs INR models to “represent” the video by learning a coordinate-based mapping and compressing the network parameters for transmission. This approach converts a video coding problem into a model compression task, which allows the use of a much smaller network to “overfit” the input video, with the real potential for fast decoding. Existing implicit neural video representation (NeRV) models can be classified into index-based and content-based methods. The former takes frame [240], patch [241] or disentangled spatial/grid coordinates [364] as model input, while content-based approaches [40, 242, 243] have content-specific embedding as inputs. Currently, one of the best INR-based video codecs [244] has already achieved a performance similar to that of VVC VTM (RA), but with a much lower decoding complexity compared to autoencoder-based neural codecs. Some of these models have also been applied to volumetric video content [246, 247], demonstrating their potential to compete with standard and other learning-based methods. However, it should be noted that the training of most NeRV models is based on an entire video sequence or even datasets; this results in a high system delay and does not meet the requirement of many low latency video streaming or real-time applications. To address this limitation, significant advances have been made [245] towards more practical INR-based video compression (such as the Low Delay and Random Access modes in VVC VTM [365]) by combining pre-training and online model overfitting.
@@
Similarly to image compression, international grand challenges are used to compare neural video compression methods, with notable venues including the NN-based Video Coding Grand Challenge associated with The IEEE International Symposium on Circuits and Systems (ISCAS) and the Challenge on Learned Image Compression (CLIC, video coding track) with IEEE/CVF CVPR and Data Compression Conference (in 2024). The best performer in ISCAS 2024 NN-based Video Coding Grand Challenge offers an overall 55% BD-rate saving over HEVC Test Model HM [366], while the winner of the CLIC (video coding track) in 2024, a neural-network enhanced ECM codec [367] with a CNN-based in-loop filter, shows a more than 2dB (in PSNR) gain compared to VTM (RA) at the same bitrates.
@@ 3.6.3 Audio Compression
Similarly to images and videos, learning-based solutions have also been researched to compress audio signals, and most neural audio codecs are based on VQ-VAE [265]. SoundStream [368] is one of such models, which can encode audio content at various bitrates. It is based on a residual vector quantizer (RVQ) which trades off between rate, distortion, and complexity. This work has been further enhanced with a multi-scale spectrogram adversary and a loss balancer mechanism, resulting in improved rate-distortion performance. A more advanced universal model has been further developed [369] based on improved adversarial and reconstruction losses, which can compress different types of audio. RVQ has also been extended from a single scale to multiple scales [370], which performs hierarchical quantization at variable frame rates.
More recently, researchers have started to exploit the use of LLMs for audio compression, leveraging the audio generation/synthesis abilities of generative models. UniAudio 1.5 [371] is one of such attempts, which converts an audio into the textural space, which can be represented by a pre-trained LLM that shares a similar backbone of UniAudio [372], a universal audio foundation model. LFSC is another neural audio codec based on LLMs, which achieved fast LLM training and inference through finite scalar quantization and adversarial training.
@@ 3.7 Visual Quality Assessment
Assessing the quality of visual signals remains an important and challenging task for many image and video processing applications. While subjective tests involving human participants remain the gold standard, objective quality models are frequently used because of their time and cost efficiency. These quality assessment methods are typically used to evaluate the performance of different visual processing approaches, and they can also be converted to loss functions, which are employed for optimizing learning-based processing models.
In recent years, quality assessment methods have been enhanced using deep learning techniques. The resulting learning-based quality models can quickly adapt to a specific type of content, leading to better performance compared to conventional, hand-crafted quality metrics. This section provides a brief summary of existing works in this research area, and highlights the main challenges which should be addressed in the near future. A more comprehensive overview of the image and video quality assessment literature can be found in [373, 374, 375].
@@ 3.7.1 Quality assessment models
Image and video quality assessment methods can be classified into two primary categories according to the availability of the corresponding reference content to the distorted test version: full-reference and no-reference models [^55] Reduced-reference quality metrics do existing in the literature, but the research in this field is less active in recent years.. Prior to the AI era, conventional visual quality methods often exploit different characteristics of the human vision system and capture relevant information related to structural similarity (such as in SSIM and its variants [376, 377, 378]), distortion [379, 380, 381], and artifacts [382, 383, 384]. In many cases, the extracted features are further processed by models that simulate texture masking [385], contrast sensitivity [386], and saliency [387]. These hand-crafted quality models have also been combined with features within a regression-based framework in order to achieve more accurate prediction performance - VMAF is one such example [335]. When neural networks are involved for feature extraction, they are trained to capture information which can directly contribute to quality prediction through an end-to-end optimization strategy. Initially, convolutional neural networks were typically used, with notable examples such as DeepQA [388], LPIPS [389] and CONTRIQUE [390] for image quality assessment, and TLVQA [391], C3DVQA [392] and DeepVQA [393] for video quality assessment. Recent works have been reported to achieve better performance when Vision Transformers (ViTs) (or similar variants) are employed due to the effectiveness of their self-attention mechanism. Important works in this class include IQT [248], TRes [249], SaTQA [250], FastVQA [251] and RankDVQA [252]. The former has been further extended as DOVER [253] and COVER [254] when aesthetic and/or semantic aspects in the content are taken into account.
@@
More recently, inspired by the success of large language models (LLMs) [4, 394] in other machine learning tasks, these have been utilized in image and video quality assessment, demonstrating significant potential to achieve better model generalization. Q-Bench [395] is one of the first attempts that employs multimodal large language models to predict the perceptual quality of images based on prompt-driven evaluation. It queries the LLMs to provide information related to the final quality rating of the input image and the quality description. This has been further extended for video quality assessment tasks in Q-Align [396]. Other notable works include X-iqe [397] that performs the quality prompt in a multi-iteration manner focusing on both image fidelity and aesthetics. Prompt-based approaches have also been proposed for differentiating the quality difference between multiple images, such as 2AFC-LMMs [398] based on a two-alternative forced choice prompt and MAP (maximum a posteriori) estimation. Moreover, recent research works also focus on using pre-trained vision-language models, such as CLIP [263], which align better image and text modalities. Important examples in this class for image quality assessment include ZEN-IQA [399], QA-CLIP [400] and PromptIQA [401]. Similar works have also been proposed for video quality assessment, such as BVQI [402, 403] and COVER [254].
@@
To support the training and validation of learning-based quality assessment models, image or video databases containing ground-truth subjective quality scores are typically employed. Commonly used image quality databases include LIVE [404], CSIQ [380], TID2013 [405], PieAPP and PIPAL, while video quality databases such as LIVE-VQA [406], KoNViD-1K [407], YouTube UGC [408] and LIVE-VQC [409] are typically employed for benchmarking in the literature. There are also databases developed that investigate the impact of specific video formats and/or artifacts, such as LIVE-YT-HFR [410] focusing on frame rates, VSR-QAD [411] on spatial resolution (or super-resolution artifacts), BAND-2k [412] on banding artifacts and Maxwell [403]/BVI-Artifact [413] containing multiple artifacts commonly produced in video streaming. Based on these databases, many learning-based quality assessment models are trained to minimize the difference (L1 or L2 norm) between predicted quality indices and subjective scores. However, due to the limited number of ground-truth quality labels associated with these databases and the resourcing-costing nature for collecting subjective data through human participants involved in psychophysical experiments, this type of training methodology cannot offer satisfactory performance, in particular when the model capacity is large. Moreover, since the experimental settings and conditions used for quality labeling are different in these databases, intra-database cross-validation is always required due to the limited model generalization and potential overfitting problems.